OpenStudy (anonymous):

I have a few math questions involving dividing radicals with extra variables... Help please?!?!?!

OpenStudy (danjs):

?

OpenStudy (anonymous):

\[3+2\sqrt{5}\div2\sqrt{5}-3\] This is one example, could anyone help explain how to work this?

OpenStudy (anonymous):

Oh That is a fraction btw...

OpenStudy (anonymous):

i am going to make a guess that this is really \[\frac{3+2\sqrt5}{2\sqrt5-3}\]

OpenStudy (danjs):

multiply top and bottom by 2*sqrt(5) +3

OpenStudy (danjs):

and simplify

OpenStudy (danjs):

remember sqrt(x)*sqrt(X) = x

OpenStudy (anonymous):

You'll need to use the difference of two squares here: \[(a+b)(a-b)=a^2+b^2\]So we multiply the top and bottom by 2sqrt{5}+3: \[\frac{3+2\sqrt{5}}{2\sqrt{5}-3}* \frac{2\sqrt{5}+3}{2\sqrt{5}+3} = \frac{(3+2\sqrt{5})(2\sqrt{5}+3)}{(2\sqrt{5})^2-3^2} \] Simplify for your answer :)

OpenStudy (danjs):

2*sqrt(5) *2*sqrt(5) = 4*5 = 20

OpenStudy (anonymous):

Okay, so my next question, how would I simplify?

OpenStudy (danjs):

3*2*sqrt(5) + 3*3 + 2sqrt(5)*2sqrt(5) + 2sqrt(5)*3 That is the top .... remember the FOIL thing

OpenStudy (danjs):

(x+y)(a+b) = x*a + x*b + y*a + y*b

OpenStudy (danjs):

I think it stood for Firsts, Outers, Inners, Lasts

OpenStudy (anonymous):

OH, Okay.

OpenStudy (danjs):

6*sqrt(5) + 9 + 4*5 + 6sqrt(5)

OpenStudy (danjs):

12 sqrt(5) + 29 that is the numerator

OpenStudy (danjs):

\[(2\sqrt{5}-3)(2\sqrt{5}+3) = \]

OpenStudy (danjs):

4*5 +6sqrt(5) - 6 sqrt(5) - 9

OpenStudy (danjs):

= 20 - 9 =11 for the Denominator

OpenStudy (anonymous):

so \[\frac{ 12\sqrt{5}+29 }{ 11 }\]

OpenStudy (danjs):

If you ever see a quantity like \[(\sqrt{x}+a)\] in the denominator, then multiply both numerator and denominator by \[(\sqrt{x}-a)\]

OpenStudy (danjs):

It will get rid of the radical in the denominator, which is not simplified

OpenStudy (anonymous):

OKay, thank you, would you be willing to help with a few more?

OpenStudy (danjs):

sure

OpenStudy (danjs):

can you open a new thread and tag me

OpenStudy (anonymous):

Alright, let me close this question and open a new one.